OpenStudy (anonymous):

Identify the 15th term of the geometric sequence 52, 39, 29.25 ... how would you set up this formula??

OpenStudy (anonymous):

same as before

OpenStudy (anonymous):

ar0+ar1...+ar14

OpenStudy (anonymous):

a is 52

OpenStudy (anonymous):

r is 39=52*r r=39/52

OpenStudy (anonymous):

52 * ((39 / 52)^14) = 0.926533297

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

i unerstand how to plug it in but not the formula.... do you have just the formula?

OpenStudy (anonymous):

the formula is just a definition i guess and is worth remembering. ar^0+ar^1+ar^2 +ar^3 notice the exponent starts at 0. And goes up by 1. THerefore the ith term is actually the the exponent -1. There is always an ar term that is worth remembering. another thing worth mentioning is that ar^0 is just a. Because the definition of anynumber^0=1 so it is actually saying a*1 which is just a. So the first term is always a.